# BELLA: Black-box model Explanations by Local Linear Approximations

### v2

@@ Abstract
Understanding the decision-making process of black-box models has become not just a legal requirement, but also an additional way to assess their performance. However, the state of the art post-hoc explanation approaches for regression models rely on synthetic data generation, which introduces uncertainty and can hurt the reliability of the explanations. Furthermore, they tend to produce explanations that apply to only very few data points. In this paper, we present BELLA, a deterministic model-agnostic post-hoc approach for explaining the individual predictions of regression black-box models. BELLA provides explanations in the form of a linear model trained in the feature space. BELLA maximizes the size of the neighborhood to which the linear model applies so that the explanations are accurate, simple, general, and robust. BELLA can produce both factual and counterfactual explanations.
@@ 1 Introduction
Numerous approaches have been proposed to build such surrogate models, in particular SHAP [Lundberg and Lee, 2017], LIME [Ribeiro et al., 2016], and MAPLE [Plumb et al., 2018]. We review them in Section 2. To evaluate the surrogate models, several criteria have been proposed: we want the surrogate model to be accurate, i.e., to reflect the predictions of the black-box model; we want it to be simple, i.e., to use few features; we want it to be robust, i.e., giving similar explanations to similar data points; we want it to be general, i.e., applicable to many data points; and we want them to be counterfactual, i.e., to tell us how we have to modify the data point at hand to get a different prediction. We survey these desiderata in Section 3. We find that existing approaches tend to be good on some of these criteria, but never excel on all of them. This is not surprising, as the desiderata stand in obvious conflict: A simple surrogate model, e.g., risks being not very accurate, because usually accurate predictions can be made only by the type of complex models that we wish to explain in the first place.
Our key idea (which we present in Section 4) is to train a local linear model on the neighborhood of the data point that we wish to explain. This allows us to develop an approach called BELLA (Black-box model Explanations by Local Linear Approximations). We can show through extensive experiments (in Section 5) on a dozen datasets that BELLA beats all existing approaches across nearly all desiderata. BELLA is freely available on Github at https://github.com/nedRad88/BELLA.
@@ 2 Related Work
Another group of approaches computes counterfactual explanations. One such method [White and Garcez, 2019] uses the idea of b-counterfactuals, i.e., the minimal change in the feature to gauge the prediction of the complex model. This method applies only to classification tasks. Another work [Dandl et al., 2020] uses Multi-Objective Optimization to compute counterfactual explanations, both for classification and regression. Another work [Redelmeier et al., 2021] uses Monte Carlo sampling for the same purpose. However, all of these approaches can compute only counterfactual explanations, not both factual and counterfactual explanations like BELLA.
@@ 3 Preliminaries
Goal. We are given a tabular dataset $T\subset F_{1}\times...\times F_{n}$, where each $F_{i}$ is a set of feature values. We are also given a function $Y:T\rightarrow\mathbb{R}$ that yields, for each $x\in T$, a target value $y\in\mathbb{R}$. These target values may, e.g., have been produced by a black-box model, in which case the target value is a prediction. Consider now one data point $x\in T$ with its target value $y$. We aim to compute an explanation in the following sense [Das and Rad, 2020]:
@@
If the target value was produced by a black-box model, we cannot be sure post-hoc that the features we identify really contributed to the computation of the target value (the model may just as well have thrown a dice, independently of any feature values). However, if several data points with these or similar feature values produce a similar prediction, we can use abductive reasoning to infer that these features may have contributed to the prediction, and that, hence, the features constitute an explanation. This is in fact common in the literature [Ribeiro et al., 2016, Lundberg and Lee, 2017, Radulovic et al., 2021, Ignatiev et al., 2019].
@@
Desiderata. In addition to the above quality measures, there are also several criteria in the literature that either apply or don’t apply to a given method of explanation. One of them is [Miller, 2018, Wachter et al., 2017]:
@@
    Counterfactuality: the ability to provide a set of modifications to the data point at hand that would entail a change in the decision of the black-box model.
Counterfactuality is obviously of interest to a user who wishes not just to understand the prediction, but also to actively influence it (e.g., after having been attributed a high risk of diabetes based on the results of a blood test).
Several methods for post-hoc explainability use randomization to probe the black-box model. However, this entails that the same data can lead to different explanations, which introduces uncertainty for the user [Zhang et al., 2019, Slack et al., 2020]. We thus have an additional desideratum:
@@
    Determinism: the avoidance of randomization steps
Finally, some methods [Plumb et al., 2018] propose explanations that take the form of a linear equation, which allows computing the predicted value from the feature values. This is a very attractive property, as the user can toy with the explanation and apply it also to neighboring data points. We thus have a desideratum that we call
  1. 7.
@@
Given this plethora of quality measures and desiderata, it is not surprising that no existing method (including our own) can satisfy all of them perfectly. However, we can at least show that our method ticks all desiderata, and outperforms existing methods across nearly all quality measures.
@@ 4 BELLA
We are given a tabular dataset $T$, a data point $x\in T$, and a real-valued target value $y$, and we aim to compute an explanation for this target value. Note that, different from other methods such as LIME and SHAP, we need as input only the dataset $T$, and not in addition the model that generated the dataset. This means that, different from SHAP and LIME, our method can explain any labeled dataset, whether it is produced by a model or comes from another source (e.g., housing prices from real estate data).
To explain the target value $y$, our idea is to find a linear equation $y\approx w_{1}\cdot x_{1}+w_{2}\cdot x_{2}+\dots+w_{l}\cdot x_{l}+w_{0}$, where $w_{i}$ are real-valued regression coefficients and $x_{i}$ are feature values of $x$ in $T$. Such an equation tells the user (1) what the important features are and (2) how they can be used to compute the predicted value. To find this equation, BELLA proceeds in three steps (Algorithm 1):
@@
3:$d\leftarrow\textrm{ComputeDistances}(x,T)$
4:$L,N\leftarrow\textrm{NeighborhoodSearch}(x,T,d)$
@@
Step 2: Neighborhood Search. After computing the distances, we proceed with the exploration of the neighborhood of the input data point $x$. The goal is to find a set of points, closest to $x$ according to the distance measure, that will serve as a training set for a local surrogate model. Several common techniques could be considered to that end, including kNN, K-Means, and other distance-based clustering methods. In our case, however, we aim to find a neighborhood such that a linear regression model trained on that neighborhood represents an accurate local approximation of the black-box model. Hence, the quality of the neighborhood is proportional to the quality of the performance of the linear model fitted on it. Common drawbacks of regression evaluation metrics are missing interpretability, sensitivity to outliers and near-zero values, divisions by zero, missing bounds, and missing symmetry. We find that the Berry-Mielke universal R value $\Re$ [Berry and Mielke Jr, 1988] avoids most of these pitfalls. $\Re$ represents the measure of agreement between raters and it is a generalization of Cohen’s kappa [Cohen, 1960]. $\Re$ measures how much better the model is compared to a naive one (e.g., to a random predictor). $\Re$ takes values from the range $[0,1]$, and it can be interpreted easily: If $\Re$ is equal to 00, the model performance is equal to the one of the random model and if it is $1$, then the model has perfect performance. $\Re$ is defined as $\Re=1-\frac{\delta}{\mu}$, where $\delta$ and $\mu$ are defined as:
@@
Here, $n$ is the number of samples, $y_{i}$ is the actual target value, $\hat{y}_{i}$ is the predicted value, and $\Delta(\cdot)$ represents the distance function between the true and the predicted value. The original work by [Berry and Mielke Jr, 1988] uses the Euclidean distance, but later works [Janson and Olsson, 2001, 2004] propose to use the squared Euclidean distance instead, because this distance is equivalent to the variance of the variable, which further improves the interpretability of $\Re$. We follow this argumentation, and use $\Delta(a,b)=(a-b)^{2}$. This definition implies that $\Delta$ is in fact equal to the Mean Squared Error (MSE). Thus, by optimizing $\Re$, we are actually optimizing the accuracy of the local model.
[…]
@@
5:$n\leftarrow$ number of features in $T$
6:$max\Re_{lb}\leftarrow 0,\textit{bestN}\leftarrow 0,\textit{bestL}\leftarrow\emptyset$
@@
1:Input: Neighborhood of data points $N$
2:$F\leftarrow$ the set of all features in $N$
3:$F^{\prime}\leftarrow\{f|f\in F\wedge\textit{VIF}(f)<10.0\}$
4:$\textit{Features}_{Lasso}\leftarrow Lasso(cv=5,\textit{features}=F^{\prime})$
@@
After removing highly collinear features, the next step is to train a linear model with Lasso regularization. Lasso regularization adds a penalty term in the form of the sum of absolute values of the regression coefficients. The objective function is $\min_{\beta\in\mathbb{R}^{p}}(||y-\beta X||_{2}^{2}+\lambda||\beta||_{1})$, where $\lambda$ is the shrinkage parameter. This provides a sparse model, by forcing some coefficients to be zero. Removing some features ensures a better generalization, and results in simpler, and thus more comprehensible explanations. On the other hand, coefficients obtained by minimizing the Lasso objective function are biased towards zero. Therefore, Lasso is preferred for model selection rather than for prediction. The common strategy is to train an Ordinary Least Squares (OLS) linear model on the subset of variables selected by Lasso. This corresponds to a special variation of the relaxed Lasso [Meinshausen, 2007], with $\phi=0$.
@@
Once we have obtained the most parsimonious model, i.e., the best set of features, we train the final local surrogate model as an OLS model using the features selected by Lasso. This procedure is described in Algorithm 3, and returns a local linear model.
This procedure is designed to maximize the robustness of our explanations: the Lasso regularization makes our results less brittle. Also, our target metric increases the number of data points, making the result more general and thus more robust.
Providing an explanation. As the final result, BELLA outputs the OLS model computed by Algorithm 3, together with the size of the neighborhood. As an example, consider the Iranian Churn dataset [Jafari-Marandi et al., 2020]. It contains the (anonymized) customers of a telecommunication company, with their age, subscription length, satisfaction with the service, etc. The goal is to predict the commercial value of the customer to the company (in dollars).
Let us now consider a given customer, for which a black-box model predicted a commercial value of \$551. The explanation that BELLA can provide for this prediction is shown in Figure 2.
@@
All numerical features have been standardized to have a mean value equal to 00 and a standard deviation equal to $1$. (Thus, a customer has a “negative age” if they are younger than the average customer.) In the explanation, the base value is the output of the model when all inputs are set to zero (i.e. to their mean value). Each bar shows the total contribution of each feature to the predicted value – positive contributions are the blue upward arrows, and negative contributions are the red downward arrows. The more the customer phones (variable seconds), the more revenue the company generates. The age (which is below average for this particular customer), likewise, has a small positive impact. The number of SMS, in contrast, (variable freqSMS) impacts the revenue negatively. Finally, the number of distinct phone numbers called (variable distnum) has a small negative impact. These sizes of the bars are easy to interpret: The size of each bar is equal to the value of the feature multiplied by the weight computed by our method. Their sum is then directly equivalent to the explained value:
@@
$\begin{split}y\approx 458.47+190.27\times\textit{seconds}-102.91\times\textit{age}\\ -480.08\times\textit{freqSMS}-17.71\times\textit{distNums}\end{split}$
@@
This computation applies to all data points in the neighborhood of the input data point (to the current instance and 476 others in our example). We thus see that BELLA’s explanations are verifiable (because they take the form of a linear equation), deterministic (because BELLA does not use any randomized steps), simple (because we applied regularization), general (because we maximized the neighborhood), and accurate (because we optimized the linear model on the local neighborhood). In addition, BELLA does not probe the black-box model. This means that, unlike many of its competitors, BELLA can explain not just the decisions of a black-box model, but any numerical variable in a tabular dataset – even if that variable was not generated by a model at all but merely observed in reality (such as, e.g., housing prices). Let us now turn to the missing desideratum, counterfactuality.
Counterfactual explanations provide information about a (minimal) change needed to alter the prediction of the black-box model. In a classification scenario, the goal is to make the model predict a different class. In a regression scenario, the goal is not to make the model predict any other value, but the value that the user would like to see. In our example of the Iranian churn dataset, an analyst may ask why the model predicted \$551 instead of, say, \$1000. A counterfactual explanation should suggest a set of changes that should be applied to reach this reference value. To provide such an explanation with BELLA, we select candidates, i.e. data points whose target value is in an $\epsilon$-vicinity to the reference value (with $\epsilon=5\%$). There can be multiple candidates.
To find the best one, we optimize two criteria: the distance between the given data point and the candidate, and the amount of change needed. The first criterion will favor candidates that are in the vicinity of the given data point. The second criterion controls the amount of change applied. To alter the outcome, one can usually either apply a small change to several features, or a big change to few features. The second approach is risky: without human intervention, we can end up with a set of features that are difficult or impossible to change (e.g., the age of a customer). Therefore, we rather aim to minimize the average amount of change and suggest smaller adjustments to multiple features (such as frequency of use, or the number of SMS messages). This yields the following objective function:
$\min_{x^{\prime}\in T}(d(x,x^{\prime})+\frac{d(x,\hat{x})}{\textit{len(Explanation)}})$ (4)
Here, $x^{\prime}$ is a counterfactual candidate data point, $d$ is a distance measure defined in Equation 1, $\hat{x}$ represents the modified data point $x$ according to the counterfactual explanation and len(Explanation) is the number of features that have been modified. The modified data point, $\hat{x}$, represents the counterfactual explanation.
The overall process is described in Algorithm 4. The algorithm takes as input the labeled dataset $T$, a labeled data point $x\in T$, a reference value $y_{\textit{ref}}\in\mathbb{R}$, and a permitted deviation $\epsilon$ from the reference value. We first choose the set of counterfactual candidates $X^{\prime}$ that have the target value in the $\epsilon$ neighborhood of the reference value $y_{\textit{ref}}$. For each $x^{\prime}$ among these candidates, we compute the explanation using BELLA [^1] In principle, any algorithm that produces a linear model can be used in Algorithm 4.. This explanation gives us a set of features, and the proposed modification of $x$ is to set all these features of $x$ to the values given by $x^{\prime}$. Among these proposed modifications, we choose the one that minimizes the objective function in Equation 4.
Such a counterfactual explanation may not always be possible: It can happen, for example, that there is no data point in the $\epsilon$ neighborhood. In such a case, the neighborhood has to be enlarged, leading to more remote example points. This is the price to pay for the fact that BELLA only ever produces real examples, and never synthetic examples. If the proposed counterfactual modifications do not lead to an actual outcome from the black-box within the desired target range, one can reduce the impact of Lasso regularization to keep even the smaller feature modifications.
1:Input: Dataset $T$ of data points $x_{i}$ with labels $y_{i}$
2:           Labeled data point $x\in T$
3:           A reference value $y_{\textit{ref}}\in\mathbb{R}$
4:           Deviation from reference value $\epsilon=0.05$
5:$X^{\prime}\leftarrow\{x_{i}\in T:y_{i}\in[y_{\textit{ref}}-\epsilon,y_{\textit{ref}}+\epsilon]\}$
6:for $x^{\prime}_{i}\in X^{\prime}$ do
7:     $L_{i},N_{i}\leftarrow\textrm{BELLA}(T,x^{\prime}_{i})$triangleright Algorithm 1
8:     $\hat{x}_{i}\leftarrow x^{\prime}_{i}$ modified according to $L_{i}$
9:end for
10:$x_{\textit{ref}}\leftarrow\textit{argmin}_{\hat{x}_{i}}(d(x,x^{\prime}_{i})+\frac{d(x,\hat{x}_{i})}{\textit{len(Explanation)}})$
11:return $x_{\textit{ref}}$
Algorithm 4 Computing a counterfactual explanation
@@
Auto MPG $7$ $6$ $1$ $392$
Bike $12$ $9$ $3$ $8760$
Concrete $8$ $8$ 00 $1030$
Servo $4$ 00 $4$ $167$
Electrical $12$ $12$ 00 $10000$
Superconductivity $81$ $81$ 00 $21262$
White Wine Quality $11$ $11$ 00 $4898$
Real Estate Valuation $5$ $5$ 00 $414$
Wind $14$ $14$ 00 $6574$
CPU activity $12$ $12$ 00 $8192$
Echocardiogram $9$ $6$ $3$ $17496$
Iranian Churn $11$ $8$ $3$ $3150$

### v3

@@ Abstract
Understanding the decision-making process of black-box models has become not just a legal requirement, but also an additional way to assess their performance. However, the state of the art post-hoc explanation approaches for regression models rely on synthetic data generation, which introduces uncertainty and can hurt the reliability of the explanations. Furthermore, they tend to produce explanations that apply to only very few data points. In this paper, we present BELLA, a deterministic model-agnostic post-hoc approach for explaining the individual predictions of regression black-box models. BELLA provides explanations in the form of a linear model trained in the feature space. BELLA maximizes the size of the neighborhood to which the linear model applies so that the explanations are accurate, simple, general, and robust.
@@ 1 Introduction
Numerous approaches have been proposed to build such surrogate models, in particular SHAP [Lundberg and Lee, 2017], LIME [Ribeiro et al., 2016], and MAPLE [Plumb et al., 2018]. We review them in Section 2. To evaluate the surrogate models, several criteria have been proposed: we want the surrogate model to be accurate, i.e., to reflect the predictions of the black-box model; we want it to be simple, i.e., to use few features; we want it to be robust, i.e., giving similar explanations to similar data points; and we want it to be general, i.e., applicable to many data points. We survey these desiderata in Section 3. We find that existing approaches tend to be good on some of these criteria, but never excel on all of them. This is not surprising, as the desiderata stand in obvious conflict: A simple surrogate model, e.g., risks being not very accurate, because usually accurate predictions can be made only by the type of complex models that we wish to explain in the first place.
Our key idea (which we present in Section 4) is to train a local linear model on the neighborhood of the data point that we wish to explain. This allows us to develop an approach called BELLA (Black-box model Explanations by Local Linear Approximations). We can show through extensive experiments (in Section 5) on a dozen datasets that BELLA beats all existing approaches across nearly all desiderata.
@@ 2 Related Work
Another group of approaches computes counterfactual explanations. One such method [White and Garcez, 2019] uses the idea of b-counterfactuals, i.e., the minimal change in the feature to gauge the prediction of the complex model. This method applies only to classification tasks. Another work [Dandl et al., 2020] uses Multi-Objective Optimization to compute counterfactual explanations, both for classification and regression. Another work [Redelmeier et al., 2021] uses Monte Carlo sampling for the same purpose. However, we focus on factual explanations in this paper.
@@ 3 Preliminaries
Goal. We are given a tabular dataset $T\subset F_{1}\times...\times F_{n}$, where each $F_{i}$ is a set of feature values. For example, a feature for a real estate dataset could be the size of the lot in square meters, and the set of feature values would then be $\mathbb{R}^{+}$. We are also given a function $Y:T\rightarrow\mathbb{R}$ that yields, for each $x\in T$, a label $Y(x)\in\mathbb{R}$. In our example, the labels can be the prices of the houses. These labels may, e.g., have been produced by a black-box model, in which case the label is a prediction. Consider now one data point $x\in T$ with its label $Y(x)$. We aim to compute an explanation in the following sense [Das and Rad, 2020]:
@@
If the label was produced by a black-box model, we cannot be sure post-hoc that the features we identify really contributed to the computation of the label (the model may just as well have thrown a dice, independently of any feature values). However, if several data points with these or similar feature values produce a similar prediction, we can use abductive reasoning to infer that these features may have contributed to the prediction, and that, hence, the features constitute an explanation. This is in fact common in the literature [Ribeiro et al., 2016, Lundberg and Lee, 2017, Radulovic et al., 2021, Ignatiev et al., 2019].
@@
Additional desiderata. In addition to the above scalar quality measures, there are also criteria in the literature that either apply or don’t apply to a given method of explanation. For example, several methods for post-hoc explainability use randomization to probe the black-box model. However, this entails that the same data can lead to different explanations, which introduces uncertainty for the user [Zhang et al., 2019, Slack et al., 2020]. We thus have as desideratum:
@@
    Determinism: the avoidance of randomization steps
Furthermore, some methods [Plumb et al., 2018] propose explanations that take the form of a linear equation, which allows computing the predicted value from the feature values. This is a very attractive property, as the user can toy with the explanation and apply it also to neighboring data points. We thus have a desideratum that we call
@@
Given this number of quality measures and desiderata, it is not surprising that no existing method (including our own) can satisfy all of them perfectly. However, we can at least show that our method ticks all desiderata, and outperforms existing methods across nearly all quality measures.
@@ 4 BELLA
We are given a tabular dataset $T$, real-valued labels $Y:T\rightarrow\mathbb{R}$, and a data point $x\in T$, and we aim to compute an explanation for the label $Y(x)$. Note that, different from other methods such as LIME and SHAP, we need as input only the dataset $T$ and the labels $Y$ on $T$, and not in addition also the model that generated the labels $Y$. This means that, different from SHAP and LIME, our method can explain any labeled dataset, whether the labels were produced by a model or whether they come from any other source (e.g., housing prices from real estate data). If BELLA is to explain a model, we first run the model on the training dataset. We then use as labels $Y$ not the ground truth from the training dataset, but the predictions of the model.
To explain the label $Y(x)$, our idea is to find a linear equation $Y(x)\approx w_{1}\cdot x_{1}+w_{2}\cdot x_{2}+\dots+w_{l}\cdot x_{l}+w_{0}$, where $w_{i}$ are real-valued regression coefficients and $x_{i}$ are feature values of $x$ in $T$. Such an equation tells the user (1) what the important features are and (2) how they can be used to compute the predicted value. To find this equation, BELLA proceeds in three steps (Algorithm 1):
@@
[…]
@@
Step 2: Neighborhood Search. After computing the distances, we proceed with the exploration of the neighborhood of the input data point $x$. The goal is to find a set of points, closest to $x$ according to the distance measure, that will serve as a training set for a local surrogate model. Several common techniques could be considered to that end, including kNN, K-Means, and other distance-based clustering methods. In our case, however, we aim to find a neighborhood such that a linear regression model trained on that neighborhood represents an accurate local approximation of the black-box model. Hence, the quality of the neighborhood is proportional to the quality of the performance of the linear model fitted on it. Common drawbacks of regression evaluation metrics are missing interpretability, sensitivity to outliers and near-zero values, divisions by zero, missing bounds, and missing symmetry. We find that the Berry-Mielke universal R value $\Re$ [Berry and Mielke Jr, 1988] avoids most of these pitfalls. $\Re$ represents the measure of agreement between raters and it is a generalization of Cohen’s kappa [Cohen, 1960]. $\Re$ measures how much better the model is compared to a naive one (e.g., to a random predictor). $\Re$ takes values from the range $[0,1]$, and it can be interpreted easily: If $\Re$ is equal to 0, the model performance is equal to the one of the random model and if it is $1$, then the model has perfect performance. $\Re$ is defined as $\Re=1-\frac{\delta}{\mu}$, where $\delta$ and $\mu$ are defined as:
@@
Here, $N$ is the number of samples, $y_{i}$ is the actual label, $\hat{y}_{i}$ is the predicted value, and $\Delta(\cdot)$ represents the distance function between the true and the predicted value. The original work by [Berry and Mielke Jr, 1988] uses the Euclidean distance, but later works [Janson and Olsson, 2001, 2004] propose to use the squared Euclidean distance instead, because this distance is equivalent to the variance of the variable, which further improves the interpretability of $\Re$. We follow this argumentation, and use $\Delta(a,b)=(a-b)^{2}$. This definition implies that $\Delta$ is in fact equal to the Mean Squared Error (MSE). Thus, by optimizing $\Re$, we are actually optimizing the accuracy of the local model.
yygIxxgI
yygIxxgI
@@
[…]
@@
The algorithm receives as input a labeled data point $x$ that is to be explained, a labeled training set $T$, and a vector of distances between $x$ and each point in the training set $T$. We sort the training set by increasing distance to $x$, train a linear model on the first $i$ data points for increasing $i$, and return the set of neighbors for which the lower bound of $\Re$ is maximal. As the neighborhood is very small in the beginning, the training easily leads to overfitting. Therefore, we consider at least 2nn data points for our neighborhood, where $n$ is the number of features. This ensures that the estimation of regression coefficients exhibits less than $10\%$ relative bias [Austin and Steyerberg, 2015].
1:Input: Neighborhood of data points $\mathcal{N}$
@@
[…]
@@
[…]
@@
Once we have obtained the most parsimonious model, i.e., the best set of features, we train the final local surrogate model as an OLS model using the features selected by Lasso. This procedure is described in Algorithm 3, and returns a local linear model. The method is designed to maximize the robustness of our explanations: the Lasso regularization makes our results less brittle. Also, our target metric increases the number of data points, making the result more general and thus more robust.
Providing an explanation. As the final result, BELLA outputs the OLS model computed by Algorithm 3, together with the size of the neighborhood. As an example, consider the Iranian Churn dataset [Jafari-Marandi et al., 2020]. It contains the (anonymized) customers of a telecommunication company, with their age, subscription length, satisfaction with the service, etc. The goal is to predict the commercial value of the customer to the company (in dollars). Let us now consider a given customer, for whom a black-box model predicted a commercial value of \$551. The explanation that BELLA can provide for this prediction is shown in Figure 2.
@@
All numerical features have been standardized to have a mean value equal to 0 and a standard deviation equal to $1$. (Thus, a customer has a “negative age” if they are younger than the average customer.) In the explanation, the base value is the output of the model when all inputs are set to zero (i.e. to their mean value). Each bar shows the total contribution of each feature to the predicted value – positive contributions are the blue upward arrows, and negative contributions are the red downward arrows. The more the customer phones (variable seconds), the more revenue the company generates. The age (which is below average for this particular customer), likewise, has a small positive impact. The number of SMS, in contrast, (variable freqSMS) impacts the revenue negatively. Finally, the number of distinct phone numbers called (variable distnum) has a small negative impact. These sizes of the bars are easy to interpret: The size of each bar is equal to the value of the feature multiplied by the weight computed by our method. Their sum is then directly equivalent to the explained value:
@@
$\begin{split}y\approx 458.47+190.27\times\textit{seconds}-102.91\times\textit{age}-480.08\times\textit{freqSMS}-17.71\times\textit{distNums}\end{split}$
@@
This computation applies to all data points in the neighborhood of the input data point (to the current instance and 476 others in our example). We thus see that BELLA’s explanations are verifiable (because they take the form of a linear equation), deterministic (because BELLA does not use any randomized steps), simple (because we applied regularization), general (because we maximized the neighborhood), and accurate (because we optimized the linear model on the local neighborhood). In addition, BELLA does not probe the black-box model. This means that, unlike many of its competitors, BELLA can explain not just the decisions of a black-box model, but any numerical variable in a tabular dataset – even if that variable was not generated by a model at all but merely observed in reality.
@@
Auto MPG 77 66 11 $392$
Bike $12$ 99 33 $8760$
Concrete 88 88 0 $1030$
Servo 44 0 44 $167$
Electrical $12$ $12$ 0 $10000$
Superconductivity $81$ $81$ 0 $21262$
White Wine Quality $11$ $11$ 0 $4898$
Real Estate Valuation 55 55 0 $414$
Wind $14$ $14$ 0 $6574$
CPU activity $12$ $12$ 0 $8192$
Echocardiogram 99 66 33 $17496$
Iranian Churn $11$ 88 33 $3150$